OpenStudy (anonymous):

Compare the values of the underlined digits. 34,258 and 47,163 - - The value of 4 in _________ is______ times the value of 4 in ________ PLEASE ANSWER ASAP IT'S FOR MY SISTER

OpenStudy (anonymous):

I believe its "The value of 4 in the thousand is \[\frac{ 1 }{ 10 }\] times the value of 4 in ten thousand. Sorry if I messed it up.

OpenStudy (anonymous):

the value of 4 in 34,258 is in the thousands place. The of 4 in 47,163 is in the 10 thousands place. comparing the two, the value of 4 in 47,163 is 10 times the value of 4 in 34,258.

OpenStudy (anonymous):

note: thousandths = 1/1000 where as thousands = 1000